OpenStudy (anonymous):

Find the distance between (–2, –3) and (–6, –12). Round the answer to the nearest hundredth, if necessary. ( answer more appreciated ) sorry i just don't get geometry

OpenStudy (austinl):

\[d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\]

OpenStudy (anonymous):

\[\sqrt{(x2 - x1)^2 + (y2 - y1)^2}\] \[\sqrt{(-6 + 2)^2 + (-12 + 3)^2} = \sqrt{( -4)^2 + (-9)^2} = 16 + 81=\sqrt{97}\]

OpenStudy (anonymous):

ANSWERS : a. 9.85 units c. 13 units b. 7.26 units d. 17.49 units

OpenStudy (anonymous):

it is A

OpenStudy (anonymous):

I just noticed that ... lol thankyou